### v1

@@ 5 Object Detection in Autonomous Vehicles
[197] YOLO-Vehicle-v1s 2024 ArXiv CSPDarknet53 Mono - - 93.50% - 45.40MB 226.00
@@
[195] YOLOv8-Lite 2024 ICCK C2f Mono 76.72% - 75.32% 4.39M - 85.00
@@
[196] ShuffYOLOX 2023 MDPI ShuffDet Mono - - 92.20% 35.45M 113.90MB -
@@
[199] improved YOLOv5 2023 Scientific Reports RepNAS Mono 92.90% 90.10% 90.10% - - 202.00
@@
[201] YOLOv10n 2024 NeurIPS CSPNet Mono 87.20% 72.10% 80.20% 27.02M 5.50MB 103.00
@@
[200] YOLOv9t 2024 ECCV GELAN Mono 82.10% 69.30% 77.50% 2.01M 4.43MB 48.00
@@
[194] YOLOv7 2023 CVPR ELAN Mono 92.30% 97.10% 95.20% - - 132.00
@@
[193] YOLOv6n 2022 ArXiv EfficientRep Mono 85.30% 67.10% 76.00% 4.23M 8.30MB 189.00
@@
[192] YOLOv5n 2020 N/A CSPDarknet53 Mono 84.90% 70.00% 78.60% 2.51M 5.04MB 127.00
@@
[191] YOLOv3-tiny 2020 ICACCS Tiny Darknet Mono 82.90% 64.80% 72.50% 1.21M 23.20MB 276.00
@@
LiDAR sensors play a pivotal role in AV perception by providing accurate 3D geometric information of the surrounding environment, enabling precise object localization and size estimation under varying lighting and weather conditions. Unlike camera-based approaches that operate in the 2D image plane, LiDAR-based methods directly process point clouds, offering richer spatial representations and improved robustness to illumination changes. Over the past decade, a wide range of 3D LiDAR-based object detection architectures have emerged, each employing different spatial representations and learning strategies to balance accuracy, efficiency, and robustness [202]. In this survey, we categorize existing approaches into six main groups, including BEV-based, point-based, voxel-based, pillar-based, range-based, and point-voxel hybrid methods, providing a detailed analysis of their core principles, representative algorithms, strengths, and limitations.
@@
[203] Fade3D 2025 T-ITS SFENet Point 94.86 89.28 86.55 90.92 82.00 77.49 - - - - - -
@@
[205] LION 2024 NeurIPS LION Voxel 91.40 89.29 87.26 89.25 80.97 78.28 - - - - - -
@@
[206] TSSTDet 2024 IEEE 3D CNN Voxel 95.80 92.11 89.23 91.84 85.47 80.65 75.13 69.38 64.31 95.16 76.24 71.62
@@
[207] SwiftPillars 2024 AAAI SPE Pillar - - - 88.07 79.83 77.50 55.34 50.21 45.93 86.06 64.92 60.75
@@
[208] PVT-SSD 2023 CVPR Sparse CNN Point-Voxel 95.23 91.63 86.43 90.66 82.29 76.85 - - - - - -
@@
[209] DSVT 2023 CVPR Transformer Voxel 92.04 88.33 87.81 87.89 80.90 78.66 - - - - - -
@@
[211] VoxelNeXt 2023 CVPR Sparse CNN Voxel 92.66 87.87 87.04 87.86 80.02 77.34 - - - - - -
@@
[212] 3D HANet 2023 TGRS VFE Point-Voxel 94.33 91.63 86.33 90.79 84.18 77.57 - - - - - -
@@
[213] TED-S 2023 AAAI TESP CNN Voxel - - - 93.05 87.91 85.81 72.38 67.81 63.54 93.09 75.77 71.20
@@
[214] RDIoU 2022 ECCV CT-stacked Voxel 94.90 89.75 84.67 90.65 82.30 77.26 - - - - - -
@@
[215] IA-SSD 2022 CVPR PointNet++ Point 93.14 89.48 84.42 88.87 80.32 75.10 46.51 39.03 35.60 78.35 61.94 55.70
@@
[216] PillarNet 2022 ECCV VGG-18/34 Pillar 93.66 87.85 86.64 89.94 79.01 77.31 - - - - - -
@@
[217] VoxSeT 2022 CVPR VoxSeT Voxel 92.70 89.07 86.29 88.53 82.06 77.46 - - - - - -
@@
[218] BtcDet 2022 AAAI RPN Voxel 92.81 89.34 84.53 90.60 83.02 77.36 69.39 61.19 55.86 91.45 74.70 70.08
@@
[219] Voxel-RCNN 2021 AAAI RPN Voxel 93.54 91.18 88.91 92.38 85.29 82.86 64.88 57.32 52.11 89.25 72.52 67.03
@@
[220] CIA-SSD 2021 AAAI SPConvNet Voxel 93.74 89.84 82.39 89.59 80.28 72.87 - - - - - -
@@
[221] CT3D 2021 ICCV RPN Voxel 92.36 88.83 84.07 87.83 81.77 77.16 65.73 58.56 53.04 91.99 71.60 67.34
@@
[223] SE-SSD 2021 CVPR SPConvNet Voxel 95.68 91.84 87.62 91.49 82.54 77.15 67.98 59.72 54.83 91.77 72.54 68.78
@@
[224] RangeDet 2021 ICCV FPN Range 87.96 69.03 48.88 - - - 82.20 75.39 65.74 - - -
@@
[226] Part-A2 2020 IEEE SPConvNet Point-Voxel 84.76 89.52 81.47 77.86 85.94 72.00 44.50 54.49 42.36 62.73 75.58 57.74
@@
[227] Point-GNN 2020 CVPR GNN Point 93.11 89.17 83.90 88.33 79.47 72.29 - - - - - -
@@
[228] 3DSSD 2020 CVPR SA Point 92.66 89.02 85.86 88.36 79.57 74.55 35.03 27.76 26.08 66.69 59.00 55.62
@@
[229] TANet 2020 AAAI Attention Voxel - - - 83.81 75.38 67.66 54.92 46.67 42.42 73.84 59.86 53.46
@@
[230] SPVCNN 2020 ECCV SPVCNN Point-Voxel - - - - - - 49.20 41.40 38.40 80.10 63.70 56.20
@@
[231] PV-RCNN 2020 CVPR 3D CNN Point-Voxel 94.98 90.65 86.14 90.25 81.43 76.82 64.26 56.67 51.91 88.88 71.95 66.78
@@
[232] PointRCNN 2019 CVPR PointNet++ Point 92.13 87.39 82.72 86.96 75.64 70.70 47.98 39.37 36.01 74.96 58.82 52.53
@@
[233] STD 2019 ICCV PointNet++ Point-Voxel 89.66 65.32 76.06 86.61 77.63 76.06 53.08 44.24 41.97 78.89 62.53 55.77
@@
[234] SECOND 2018 MDPI Sparse CNN Voxel 88.07 79.37 77.95 83.13 73.66 66.20 45.31 35.52 33.14 75.83 60.82 53.67
@@
[235] VoxelNet 2018 CVPR VFE Voxel 89.60 84.81 87.57 81.97 65.46 62.85 65.95 61.05 56.98 74.41 52.18 50.49
@@
From Table 18, we have selected four models for detailed discussion, focusing on their distinctive contributions for 3D object detection in KITTI dataset: TED-S [213], obtained the highest results in $AP_{3D}$ CarCar evaluations; TSSTDet [206], which achieved the highest results in $AP_{BEV}$ CarCar and $AP_{3D}$ CyclistCyclist evaluations; SECOND [234], noted as one of the most widely adopted models in this field; and Fade3D [203], representing the latest advancement in this domain. It should be mentioned that the values highlighted in red represent the best performance achieved for each evaluation metric across all listed methods.
@@
TED-S [213] is a two-stage LiDAR 3D object detector that introduces transformation-equivariant processing into the Voxel-RCNN framework. The network first applies a Transformation-equivariant Sparse Convolution (TeSpConv) backbone, in which the input point cloud is transformed into multiple discrete yaw-rotated and mirrored copies, voxelized, and processed with shared sparse convolutional weights to produce multi-channel equivariant voxel features. These are aggregated at the scene level through TeBEV pooling, which collapses each transformation channel into a BEV map, aligns them via bilinear sampling, and max-pools across channels to form a compact, transformation-consistent representation for the 2D RPN proposal stage. Proposals are then refined using TiVoxel pooling, where instance-level grids are projected into each transformation channel, voxel features are sampled via Voxel Set Abstraction, concatenated across channels, and fused using cross-grid attention for the final detection head. This design preserves geometric consistency under predefined rigid transformations while leveraging an efficient RPN-refinement pipeline for high-accuracy 3D detection.
@@
TSSTDet framework [206] utilizes a multistage design that addresses two challenges in LiDAR-based object detection: the considerable variation in object orientation and occlusion affecting geometry. In the first stage, a Rotational-Transformation Convolutional backbone (RTConv) processes multiple rotated and reflected variants of the input point cloud through a shared sparse convolutional network, producing sets of voxel features that remain equivariant to the applied transformations. These features are aligned and merged into a unified bird’s-eye-view representation via a custom RT BEV pooling process, which feeds a region proposal network to generate initial 3D candidate boxes. Building on these proposals, the second stage introduces the Voxel-Point Shape Transformer (VPST). This autoregressive transformer module reconstructs the object geometry by predicting a sequence of quantized voxel features from the partial observation, thereby enriching the proposal with occlusion-compensated detail. The final stage uses a Fusion and Refinement network, which combines the original rotation equivariant features with the completed shape features for enabling high-confidence bounding box predictions.
@@
SECOND [234] extends voxel-based 3D detection by introducing a sparse convolutional backbone that dramatically reduces the computational footprint of earlier dense 3D convolution designs. The network begins by grouping points into voxels and applying a Voxel Feature Encoding (VFE) module to produce learned descriptors for each occupied cell. Those features enter the Sparse Convolutional Middle Extractor, a series of 3D sparse convolutional layers that operate on non-empty locations to maintain spatial fidelity without unnecessary computation in space. The vertical resolution is reduced as processing progresses, enabling the feature map to be reshaped into a dense BEV representation suitable for efficient 2D convolutional processing. This representation is then handled by a region proposal network with an SSD-like multi-scale design, combining standard and transposed convolutions to generate proposals across object sizes. This design also introduces refinements such as a sine-error formulation for orientation regression, which further enhances geometric precision and detection stability while maintaining the backbone’s streamlined nature.
@@
Fade3D framework [203] aims to minimize computational latency while maintaining the essential spatial cues for 3D object detection. The processing begins with a Lightweight Input Encoder (LIE), which converts the raw point cloud into a compact Birds Eye View (BEV) image. Rather than employing computationally intensive point-based encoders, the system relies on a pilot study showing that recording only the maximum height and maximum reflectance per grid cell achieves an optimal balance between accuracy and runtime. This encoded BEV representation is passed to SFENet, a specialized 2D convolutional backbone constructed in three stages: an initial convolutional block for base feature extraction, a depth-wise shuffle block augmented with channel attention for efficient mid-level processing, and a re-parameterized block that simplifies to a single convolution at inference time to accelerate deployment. The final stage integrates an IoU aware loss re-weighting strategy toward more challenging samples without impacting runtime, which ensures an uncompromised inference speed.
@@
[236] Voxel Mamba 2024 NeurIPS SSMs Voxel 73.0 69.0 86.8 57.1 68.0 63.2 89.5 74.7 50.8
@@
[205] LION-RetNet 2024 NeurIPS LION3D Point 71.9 67.3 89.6 64.3 78.7 44.6 89.8 73.5 56.6
@@
[205] LION-RWKV 2024 NeurIPS LION3D Voxel 71.7 67.6 89.4 59.0 77.6 37.1 89.7 74.3 56.2
@@
[205] LION-Mamba 2024 NeurIPS LION3D Voxel 72.1 68.0 87.9 64.9 77.6 44.4 89.6 75.6 59.4
@@
[237] PillarNeSt-Base 2024 T-IV ConNeXt Pillar 71.3 66.9 87.4 56.4 64.0 63.0 86.6 69.4 46.8
@@
[237] PillarNeSt-Large 2024 T-IV ConvNeXt Pillar 71.6 66.9 87.4 56.4 64.0 63.0 86.6 69.4 51.5
@@
[238] SAFDNet 2024 CVPR Sparse CNN Voxel 72.3 68.3 87.3 57.3 68.0 61.7 89.0 71.1 44.8
@@
[239] FSDv2 2024 TPAMI SparseUNet Voxel 71.7 66.5 86.1 53.0 66.5 61.1 5 87.1 71.1 51.7
@@
[211] VoxelNeXt 2023 CVPR Sparse CNN Voxel 70.0 64.5 84.6 53.0 64.7 55.8 85.8 73.2 45.7
@@
[240] LinK 2023 CVPR LinK-based Voxel 71.0 66.3 86.7 55.7 65.7 62.1 85.5 73.5 47.5
@@
[241] HEDNet 2023 NeurIPS Sparse CNN Point-Voxel 72.0 67.7 86.5 57.9 70.4 61.2 87.9 70.4 46.9
@@
[242] LargeKernel3D 2023 CVPR 3D CNN Voxel 70.9 65.4 85.5 53.8 64.4 59.5 85.9 72.7 46.8
@@
[209] DSVT 2023 CVPR BEV Voxel 72.7 68.0 86.8 54.8 67.4 63.1 88.0 73.0 47.2
@@
[243] FSTR 2023 TGRS VoxelNeXt Point 71.5 67.2 86.5 54.1 66.4 58.4 88.6 73.7 48.1
@@
[244] Uni3DETR 2023 NeurIPS Sparse CNN Point-Voxel 68.5 57.3 86.1 57.8 63.5 38.2 88.7 74.6 42.2
@@
[245] FCOS-LiDAR 2022 NeurIPS LiDAR-Net Range 65.7 60.2 82.2 47.7 52.9 48.8 84.5 68.0 39.0
@@
[246] AFDetV2 2022 AAAI RPN Voxel 68.5 62.2 86.3 52.6 62.8 59.9 85.4 68.3 34.8
@@
[186] UVTR-L 2022 NeurIPS ResNet-101 Voxel 69.7 63.9 86.3 52.2 62.8 59.7 84.5 68.8 41.1
@@
[247] Focals Conv 2022 CVPR Sparse CNN Voxel 70.0 63.8 86.7 56.3 67.7 59.5 87.5 64.5 36.3
@@
[216] PillarNet 2022 ECCV VGG-18/34 Pillar 71.4 66.4 86.1 53.1 65.3 63.1 87.3 70.1 42.3

### v2

@@ 5 Object Detection in Autonomous Vehicles
[li2024yolo] YOLO-Vehicle-v1s 2024 ArXiv CSPDarknet53 Mono - - 93.50% - 45.40MB 226.00
@@
[yang2024yolov8] YOLOv8-Lite 2024 ICCK C2f Mono 76.72% - 75.32% 4.39M - 85.00
@@
[he2023object] ShuffYOLOX 2023 MDPI ShuffDet Mono - - 92.20% 35.45M 113.90MB -
@@
[jia2023fast] improved YOLOv5 2023 Scientific Reports RepNAS Mono 92.90% 90.10% 90.10% - - 202.00
@@
[wang2024yolov10] YOLOv10n 2024 NeurIPS CSPNet Mono 87.20% 72.10% 80.20% 27.02M 5.50MB 103.00
@@
[wang2024yolov9] YOLOv9t 2024 ECCV GELAN Mono 82.10% 69.30% 77.50% 2.01M 4.43MB 48.00
@@
[wang2023yolov7] YOLOv7 2023 CVPR ELAN Mono 92.30% 97.10% 95.20% - - 132.00
@@
[li2022yolov6] YOLOv6n 2022 ArXiv EfficientRep Mono 85.30% 67.10% 76.00% 4.23M 8.30MB 189.00
@@
[gao2025algorithm] YOLOv5n 2020 N/A CSPDarknet53 Mono 84.90% 70.00% 78.60% 2.51M 5.04MB 127.00
@@
[adarsh2020yolo] YOLOv3-tiny 2020 ICACCS Tiny Darknet Mono 82.90% 64.80% 72.50% 1.21M 23.20MB 276.00
@@
LiDAR sensors play a pivotal role in AV perception by providing accurate 3D geometric information of the surrounding environment, enabling precise object localization and size estimation under varying lighting and weather conditions. Unlike camera-based approaches that operate in the 2D image plane, LiDAR-based methods directly process point clouds, offering richer spatial representations and improved robustness to illumination changes. Over the past decade, a wide range of 3D LiDAR-based object detection architectures have emerged, each employing different spatial representations and learning strategies to balance accuracy, efficiency, and robustness [mehrabi2026turbo]. In this survey, we categorize existing approaches into six main groups, including BEV-based, point-based, voxel-based, pillar-based, range-based, and point-voxel hybrid methods, providing a detailed analysis of their core principles, representative algorithms, strengths, and limitations.
@@
[ye2025fade3d] Fade3D 2025 T-ITS SFENet Point 94.86 89.28 86.55 90.92 82.00 77.49 - - - - - -
@@
[liu2024lion] LION 2024 NeurIPS LION Voxel 91.40 89.29 87.26 89.25 80.97 78.28 - - - - - -
@@
[hoang2024tsstdet] TSSTDet 2024 IEEE 3D CNN Voxel 95.80 92.11 89.23 91.84 85.47 80.65 75.13 69.38 64.31 95.16 76.24 71.62
@@
[jin2024swiftpillars] SwiftPillars 2024 AAAI SPE Pillar - - - 88.07 79.83 77.50 55.34 50.21 45.93 86.06 64.92 60.75
@@
[yang2023pvt] PVT-SSD 2023 CVPR Sparse CNN Point-Voxel 95.23 91.63 86.43 90.66 82.29 76.85 - - - - - -
@@
[wang2023dsvt] DSVT 2023 CVPR Transformer Voxel 92.04 88.33 87.81 87.89 80.90 78.66 - - - - - -
@@
[chen2023voxelnext] VoxelNeXt 2023 CVPR Sparse CNN Voxel 92.66 87.87 87.04 87.86 80.02 77.34 - - - - - -
@@
[xia20233] 3D HANet 2023 TGRS VFE Point-Voxel 94.33 91.63 86.33 90.79 84.18 77.57 - - - - - -
@@
[wu2023transformation] TED-S 2023 AAAI TESP CNN Voxel - - - 93.05 87.91 85.81 72.38 67.81 63.54 93.09 75.77 71.20
@@
[sheng2022rethinking] RDIoU 2022 ECCV CT-stacked Voxel 94.90 89.75 84.67 90.65 82.30 77.26 - - - - - -
@@
[zhang2022not] IA-SSD 2022 CVPR PointNet++ Point 93.14 89.48 84.42 88.87 80.32 75.10 46.51 39.03 35.60 78.35 61.94 55.70
@@
[shi2022pillarnet] PillarNet 2022 ECCV VGG-18/34 Pillar 93.66 87.85 86.64 89.94 79.01 77.31 - - - - - -
@@
[he2022voxel] VoxSeT 2022 CVPR VoxSeT Voxel 92.70 89.07 86.29 88.53 82.06 77.46 - - - - - -
@@
[xu2022behind] BtcDet 2022 AAAI RPN Voxel 92.81 89.34 84.53 90.60 83.02 77.36 69.39 61.19 55.86 91.45 74.70 70.08
@@
[deng2021voxel] Voxel-RCNN 2021 AAAI RPN Voxel 93.54 91.18 88.91 92.38 85.29 82.86 64.88 57.32 52.11 89.25 72.52 67.03
@@
[zheng2021cia] CIA-SSD 2021 AAAI SPConvNet Voxel 93.74 89.84 82.39 89.59 80.28 72.87 - - - - - -
@@
[sheng2021improving] CT3D 2021 ICCV RPN Voxel 92.36 88.83 84.07 87.83 81.77 77.16 65.73 58.56 53.04 91.99 71.60 67.34
@@
[zheng2021se] SE-SSD 2021 CVPR SPConvNet Voxel 95.68 91.84 87.62 91.49 82.54 77.15 67.98 59.72 54.83 91.77 72.54 68.78
@@
[fan2021rangedet] RangeDet 2021 ICCV FPN Range 87.96 69.03 48.88 - - - 82.20 75.39 65.74 - - -
@@
[shi2020points] Part-A2 2020 IEEE SPConvNet Point-Voxel 84.76 89.52 81.47 77.86 85.94 72.00 44.50 54.49 42.36 62.73 75.58 57.74
@@
[shi2020point] Point-GNN 2020 CVPR GNN Point 93.11 89.17 83.90 88.33 79.47 72.29 - - - - - -
@@
[yang20203dssd] 3DSSD 2020 CVPR SA Point 92.66 89.02 85.86 88.36 79.57 74.55 35.03 27.76 26.08 66.69 59.00 55.62
@@
[liu2020tanet] TANet 2020 AAAI Attention Voxel - - - 83.81 75.38 67.66 54.92 46.67 42.42 73.84 59.86 53.46
@@
[tang2020searching] SPVCNN 2020 ECCV SPVCNN Point-Voxel - - - - - - 49.20 41.40 38.40 80.10 63.70 56.20
@@
[shi2020pv] PV-RCNN 2020 CVPR 3D CNN Point-Voxel 94.98 90.65 86.14 90.25 81.43 76.82 64.26 56.67 51.91 88.88 71.95 66.78
@@
[shi2019pointrcnn] PointRCNN 2019 CVPR PointNet++ Point 92.13 87.39 82.72 86.96 75.64 70.70 47.98 39.37 36.01 74.96 58.82 52.53
@@
[yang2019std] STD 2019 ICCV PointNet++ Point-Voxel 89.66 65.32 76.06 86.61 77.63 76.06 53.08 44.24 41.97 78.89 62.53 55.77
@@
[yan2018second] SECOND 2018 MDPI Sparse CNN Voxel 88.07 79.37 77.95 83.13 73.66 66.20 45.31 35.52 33.14 75.83 60.82 53.67
@@
[zhou2018voxelnet] VoxelNet 2018 CVPR VFE Voxel 89.60 84.81 87.57 81.97 65.46 62.85 65.95 61.05 56.98 74.41 52.18 50.49
@@
From Table 18, we have selected four models for detailed discussion, focusing on their distinctive contributions for 3D object detection in KITTI dataset: TED-S [wu2023transformation], obtained the highest results in $AP_{3D}$ CarCar evaluations; TSSTDet [hoang2024tsstdet], which achieved the highest results in $AP_{BEV}$ CarCar and $AP_{3D}$ CyclistCyclist evaluations; SECOND [yan2018second], noted as one of the most widely adopted models in this field; and Fade3D [ye2025fade3d], representing the latest advancement in this domain. It should be mentioned that the values highlighted in red represent the best performance achieved for each evaluation metric across all listed methods.
@@
TED-S [wu2023transformation] is a two-stage LiDAR 3D object detector that introduces transformation-equivariant processing into the Voxel-RCNN framework. The network first applies a Transformation-equivariant Sparse Convolution (TeSpConv) backbone, in which the input point cloud is transformed into multiple discrete yaw-rotated and mirrored copies, voxelized, and processed with shared sparse convolutional weights to produce multi-channel equivariant voxel features. These are aggregated at the scene level through TeBEV pooling, which collapses each transformation channel into a BEV map, aligns them via bilinear sampling, and max-pools across channels to form a compact, transformation-consistent representation for the 2D RPN proposal stage. Proposals are then refined using TiVoxel pooling, where instance-level grids are projected into each transformation channel, voxel features are sampled via Voxel Set Abstraction, concatenated across channels, and fused using cross-grid attention for the final detection head. This design preserves geometric consistency under predefined rigid transformations while leveraging an efficient RPN-refinement pipeline for high-accuracy 3D detection.
@@
TSSTDet framework [hoang2024tsstdet] utilizes a multistage design that addresses two challenges in LiDAR-based object detection: the considerable variation in object orientation and occlusion affecting geometry. In the first stage, a Rotational-Transformation Convolutional backbone (RTConv) processes multiple rotated and reflected variants of the input point cloud through a shared sparse convolutional network, producing sets of voxel features that remain equivariant to the applied transformations. These features are aligned and merged into a unified bird’s-eye-view representation via a custom RT BEV pooling process, which feeds a region proposal network to generate initial 3D candidate boxes. Building on these proposals, the second stage introduces the Voxel-Point Shape Transformer (VPST). This autoregressive transformer module reconstructs the object geometry by predicting a sequence of quantized voxel features from the partial observation, thereby enriching the proposal with occlusion-compensated detail. The final stage uses a Fusion and Refinement network, which combines the original rotation equivariant features with the completed shape features for enabling high-confidence bounding box predictions.
@@
SECOND [yan2018second] extends voxel-based 3D detection by introducing a sparse convolutional backbone that dramatically reduces the computational footprint of earlier dense 3D convolution designs. The network begins by grouping points into voxels and applying a Voxel Feature Encoding (VFE) module to produce learned descriptors for each occupied cell. Those features enter the Sparse Convolutional Middle Extractor, a series of 3D sparse convolutional layers that operate on non-empty locations to maintain spatial fidelity without unnecessary computation in space. The vertical resolution is reduced as processing progresses, enabling the feature map to be reshaped into a dense BEV representation suitable for efficient 2D convolutional processing. This representation is then handled by a region proposal network with an SSD-like multi-scale design, combining standard and transposed convolutions to generate proposals across object sizes. This design also introduces refinements such as a sine-error formulation for orientation regression, which further enhances geometric precision and detection stability while maintaining the backbone’s streamlined nature.
@@
Fade3D framework [ye2025fade3d] aims to minimize computational latency while maintaining the essential spatial cues for 3D object detection. The processing begins with a Lightweight Input Encoder (LIE), which converts the raw point cloud into a compact Birds Eye View (BEV) image. Rather than employing computationally intensive point-based encoders, the system relies on a pilot study showing that recording only the maximum height and maximum reflectance per grid cell achieves an optimal balance between accuracy and runtime. This encoded BEV representation is passed to SFENet, a specialized 2D convolutional backbone constructed in three stages: an initial convolutional block for base feature extraction, a depth-wise shuffle block augmented with channel attention for efficient mid-level processing, and a re-parameterized block that simplifies to a single convolution at inference time to accelerate deployment. The final stage integrates an IoU aware loss re-weighting strategy toward more challenging samples without impacting runtime, which ensures an uncompromised inference speed.
@@
[zhang2024voxel] Voxel Mamba 2024 NeurIPS SSMs Voxel 73.0 69.0 86.8 57.1 68.0 63.2 89.5 74.7 50.8
@@
[liu2024lion] LION-RetNet 2024 NeurIPS LION3D Point 71.9 67.3 89.6 64.3 78.7 44.6 89.8 73.5 56.6
@@
[liu2024lion] LION-RWKV 2024 NeurIPS LION3D Voxel 71.7 67.6 89.4 59.0 77.6 37.1 89.7 74.3 56.2
@@
[liu2024lion] LION-Mamba 2024 NeurIPS LION3D Voxel 72.1 68.0 87.9 64.9 77.6 44.4 89.6 75.6 59.4
@@
[mao2024pillarnest] PillarNeSt-Base 2024 T-IV ConNeXt Pillar 71.3 66.9 87.4 56.4 64.0 63.0 86.6 69.4 46.8
@@
[mao2024pillarnest] PillarNeSt-Large 2024 T-IV ConvNeXt Pillar 71.6 66.9 87.4 56.4 64.0 63.0 86.6 69.4 51.5
@@
[zhang2024safdnet] SAFDNet 2024 CVPR Sparse CNN Voxel 72.3 68.3 87.3 57.3 68.0 61.7 89.0 71.1 44.8
@@
[fan2024fsd] FSDv2 2024 TPAMI SparseUNet Voxel 71.7 66.5 86.1 53.0 66.5 61.1 5 87.1 71.1 51.7
@@
[chen2023voxelnext] VoxelNeXt 2023 CVPR Sparse CNN Voxel 70.0 64.5 84.6 53.0 64.7 55.8 85.8 73.2 45.7
@@
[lu2023link] LinK 2023 CVPR LinK-based Voxel 71.0 66.3 86.7 55.7 65.7 62.1 85.5 73.5 47.5
@@
[zhang2023hednet] HEDNet 2023 NeurIPS Sparse CNN Point-Voxel 72.0 67.7 86.5 57.9 70.4 61.2 87.9 70.4 46.9
@@
[chen2023largekernel3d] LargeKernel3D 2023 CVPR 3D CNN Voxel 70.9 65.4 85.5 53.8 64.4 59.5 85.9 72.7 46.8
@@
[wang2023dsvt] DSVT 2023 CVPR BEV Voxel 72.7 68.0 86.8 54.8 67.4 63.1 88.0 73.0 47.2
@@
[zhang2023fully] FSTR 2023 TGRS VoxelNeXt Point 71.5 67.2 86.5 54.1 66.4 58.4 88.6 73.7 48.1
@@
[wang2023uni3detr] Uni3DETR 2023 NeurIPS Sparse CNN Point-Voxel 68.5 57.3 86.1 57.8 63.5 38.2 88.7 74.6 42.2
@@
[tian2022fully] FCOS-LiDAR 2022 NeurIPS LiDAR-Net Range 65.7 60.2 82.2 47.7 52.9 48.8 84.5 68.0 39.0
@@
[hu2022afdetv2] AFDetV2 2022 AAAI RPN Voxel 68.5 62.2 86.3 52.6 62.8 59.9 85.4 68.3 34.8
@@
[li2022unifying] UVTR-L 2022 NeurIPS ResNet-101 Voxel 69.7 63.9 86.3 52.2 62.8 59.7 84.5 68.8 41.1
@@
[chen2022focal] Focals Conv 2022 CVPR Sparse CNN Voxel 70.0 63.8 86.7 56.3 67.7 59.5 87.5 64.5 36.3
@@
[shi2022pillarnet] PillarNet 2022 ECCV VGG-18/34 Pillar 71.4 66.4 86.1 53.1 65.3 63.1 87.3 70.1 42.3